. Big plane. A six-passenger plane cruises at 180 mph in calm air. If the plane flies 7 miles with the wind I the same amount as it flies 5 miles against the wind, then what is the wind speed?
Answer by checkley71(8403)

7/(180+X)=5/(180-X)
5(180+X)=7(180-X)
900+5X=1260-7X
5X+7X=1260-900
12X=360
X=360/12
X=30 MPH IS THE WIND SPEED
PROOF
7/(180+30)=5/(180-30)
7/210=5/150
5*210=7*150
1050=1050